OpenStudy (anonymous):

ANSWER IN 5 MINUTES FOR MEDAL + FAN!! Call me dumb but I need help for this one :D Solve. |- x - 67| = 32

OpenStudy (anonymous):

the answer is -35

OpenStudy (igreen):

You separate this into two different equations. -x - 67 = 32 and x + 67 = 32 ---------------- -x - 67 = 32 Add 67 to both sides: -x = 99 Multiply -1 to both sides: x = -99 ------------------ x + 67 = 32 Subtract 67 to both sides: x = -35

OpenStudy (igreen):

The answer is x = -99 AND -35.

OpenStudy (anonymous):

can 1 out of 6 people help me?

OpenStudy (anonymous):

maybe @TheAsker2002

OpenStudy (anonymous):

ya @TheAsker2002

OpenStudy (anonymous):

Thank again @iGreen I'll continue trying to fan you :)

OpenStudy (anonymous):

ok @Datgirl14